25x^3 = 64x
25x^3 - 64x = 0
Not sure what to do after I set right side to 0

Found 2 solutions by John10, richard1234:
Answer by John10(297)

25x^3 = 64x
25x^3 - 64x = 0
You need to factor out the common factor
x(25x^2 -64) =0
Apply the zero product property
x = 0 OR 25x^2 -64 =0
x = 0 OR 25x^2 = 64
x= 0 OR x^2 = 64/25
x =0 OR x = +/- SR(64/25)
x =0 OR x = 8/5 OR x = -8/5

Answer by richard1234(7193)

Factor out x to obtain

x(25x^2 - 64) = 0

Either x = 0 or 25x^2 - 64 = 0. I think you can take it from there.